Two charges are found to be 0.25 m apart. The charge of the first is thrice the charge of the second. If the magnitude of the second charge is 2 uC, find the force experienced by the first charge due to the second charge.


1
Expert's answer
2022-02-16T10:23:59-0500

Given:

"r=\rm 0.25\: m"

"q_1=3q_2=6*10^{-6}\:\rm C"

"q_2=2*10^{-6}\:\rm C"


The Coulomb's law says

"F=k\frac{q_1q_2}{r^2}\\
=9*10^9*\frac{6*10^{-6}*1*10^{-6}}{(0.25)^2}=0.864\: \rm N"